MONT 106N – Identifying Patterns Discussion Day 1 September 18, 2009

Paige Carr and Andrew Coury

In class on September 18, 2009, you will be responsible for about 20 minutes of the class period, during which we will discuss the article “Facebook fans do worse in exams” from the London Times (UK), published on April 12, 2009. You can assume everyone will read the article before class – the information how to get it is posted on the course homepage. So you don’t need to spend a lot of time summarizing what it says. Instead, come up with two or three good questions that you can use to start the discussion, and then be prepared to “moderate” (maybe answer some questions if people have not understood some point, call on people who have something to say, and so forth).

I found this by looking at a listing in Chance News 48 – see link to this on the course homepage. You will see a discussion of the issues in Chance News and several other related articles that may give you a few ideas of how to get a discussion started. And of course, you can talk things over with me before class too.